A Complete Bibliography of the Journal of the Royal Statistical Society, Series A family: 1830–1839

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 April 2024 Version 1.03

Title word cross-reference

&c [141].

15th [106]. 1828 [106]. 1834 [2, 1]. 1835 [6, 10, 4, 3, 8, 5, 7, 9]. 1836 [14, 18, 12, 16, 13, 15, 17, 11]. 1837 [19, 22, 23, 27, 20, 25, 21, 24, 26, 83]. 1838 [28, 83, 110]. 1839 [147, 149, 136].

2nd [83].

31st [110].

60 [106].

Abstract [64, 97, 70, 83, 139]. Abstracted [126]. Account [64, 42, 106, 78, 77, 127, 131, 87, 35, 126, 137, 95, 110, 138, 94]. Act [106, 132]. Actual [57]. Address [80]. Advancement [82, 152, 147]. Advantages [80]. Africa [127]. Agricultural [93, 62, 43]. Agriculture [93, 148]. Algeria [127]. Among [139, 54, 69, 141, 136, 155, 99]. Amongst [76]. Amount [84]. Amputation [45]. Anne [103]. Annual [48, 31, 90, 130, 142, 125]. Antwerp [49]. Appendix [78]. Applied [121]. Appointed [68, 102, 152, 136]. Approximate [113]. April [23, 6, 14]. Army [76]. Association [82, 152, 147]. Asylum [88]. August [83, 89, 147]. Average [106].

Banks [111]. Bedford [43]. Beginning [33]. being [78]. Belgium [122]. Bellingham [96]. Belong [89]. Birth [61]. Births [142, 141, 126]. Blind [88]. Borough [137, 34]. Branches [57]. Bristol [42, 154]. Britain [140, 92]. British [106, 82, 152, 147]. Bromwich [155]. Brussels [49]. Bury [112].

c [106]. Cadiz [85]. Causes [61]. Census [144]. Central [36]. Century [57, 33]. Changes [87, 153]. Chapels [89]. Chaplain [125]. Charities [56, 138]. Charity [78]. Chaussées [84]. Cheshire [95]. Chiefly [70]. Children [32, 71]. Churches [89]. Circumstances [61]. City [42, 141, 154, 120, 103, 114]. Class [59]. Classes [155, 95, 154]. Clay [125]. Club [94]. Cognizance [83]. Collect [136]. Collection [148]. Collieries [152, 94]. Colonial [106]. Coming [83]. Commercial [131]. Commissioners [74]. Committee [68, 102, 152, 136, 124, 150, 103]. Companies [41]. Compared [109, 126]. Comparison [57]. Compiler [70]. Condition [61, 149, 87, 154, 93]. conducted [34]. Connected [61, 49]. Consumption [140, 55]. Containing [89, 137]. Conveyance [49]. Copper [40]. Corn [104]. Cornwall [56, 140, 137, 40]. Coronations [64]. Correction [125]. Council [31, 130]. County [32, 150, 96, 43]. Crime [71, 98, 151]. Criminal [61, 86, 70].

Darton [94]. Date [106]. Deaths [142, 126]. December [27, 2, 10, 18, 126]. Decennial [109]. Denmark [144]. Denominations [89]. Dependent [71]. Detailed [78]. Different [57]. Direction [84]. Diseases [55]. Distinguishing [89]. District [115, 32]. Division [93]. Dr. [90, 141]. Drawn [84]. Dundee [110]. During [72, 63, 111, 57, 83, 104, 157, 133, 153, 110]. Duty [106].

each [106, 78]. Earnings [62]. Economical [33]. Editor [100]. Education [78, 61, 68, 102, 124, 150, 155, 36, 103]. Educational [86, 132]. Effects [132]. Eighth [82]. Emigration [58]. Ended [110]. Endowed [78, 56, 138]. Engine [77]. England [142, 116, 151, 70]. Enquire [68, 102, 136]. Enquiries [97, 80]. Enquiry [42, 34, 75]. Especially [121]. Establishment [77, 32]. Excess [141]. Expenditure [33]. Expenses [64]. Explanatory [80]. Extent [131, 113].

Factorie [132]. Factories [132]. Families [42, 34]. Favour [148]. February [20, 4, 12]. Fever [75]. Fifth [130]. Fire [77]. Fire-Engine [77]. Fires [77]. First [142, 152]. Flannel [46]. Flour [106]. Force [136]. Foreign [106, 141]. Forms [61]. Fourth [64, 31]. France [41, 109, 111, 84]. French [127]. Frequency [55]. Further [72, 157].

G. [33]. General [126, 142]. Générale [84]. Geo [106]. George [64, 103]. Glasgow [59]. Glass [113]. Goods [131]. Gray [115]. Great [140, 92].

Hamburg [141]. Heart [55].

Herefordshire [138]. House [125, 34].

Houses [72, 63, 157]. Hyde [95].

III [33]. Illustrations [92]. Illustrative [71]. Immediate [79]. Imported [106]. Incidental [126]. Including [126]. Increase [79]. Increased [57]. Indian [76]. Indies [54, 69, 99]. Industry [71, 84]. Infant [78]. Information [29, 108, 119]. Inhabitants [89]. Inn [115]. Inquire [152]. Inquiry [154, 151]. Inspectors [132]. Interchange [33]. Introduction [30]. Ireland [92, 98]. Irish [74]. IV [106, 33].

James [103]. January [28, 3, 19, 11]. Jews [141]. John [78]. Joint [41]. Joint-Stock [41]. Julius [141]. July [106]. June [25, 8, 16].

Kingdom [58, 37, 50, 113, 126, 92, 139].

Labourers [93, 62]. Labouring [95]. Lancashire [79, 112]. Last [90, 144]. Letter [90, 141]. List [72, 63, 157]. Literary [156]. Liverpool [133]. London [78, 31, 77, 68, 102, 130, 141, 133, 136, 103, 115, 156].

Madron [137]. Maintained [32]. Major [139]. Male [141]. Manchester [124, 150, 34]. Manufacture [113]. Manufactured [131]. Manufactures [131]. March [21, 5, 13, 22, 149]. Margaret [78]. Marriages [142, 126]. Massachusetts [60]. May [24, 110, 7, 15, 136]. Medicine [121]. Meeting [82, 147]. Member [141]. Method [121]. Metropolis [44]. Metropolitan [136]. Miles [34]. Mines [140, 40, 84]. Mining [84]. Miscellaneaous [39, 91]. Miscellaneous [53, 67, 73, 81, 101, 107, 118, 129, 135, 146, 158]. Monday [23, 27, 20, 25, 21, 24, 26, 28]. Month [106, 89]. Monthly [106]. Months [83]. Moral [103, 115, 114]. Mortality [136, 76, 139, 45, 54, 69, 99].

Nation [33]. Nature [61, 133]. Near [115, 112]. Newcastle [89, 152, 86, 88, 83]. Newcastle-upon-Tyne [89, 86, 88, 83]. Nineteenth [33]. Ninth [147]. Norfolk [62]. Northern [93]. Northumberland [96, 93]. Norway [143]. Norwich [114]. Note [100]. Notes [71]. Notice [144, 88]. November [26, 1, 9, 17]. Number [89, 133]. Numerical [121].

Objects [80]. Observations [59]. October [83]. Offences [61]. Offenders [61]. Officers [76]. Official [84]. Operations [111]. Operatives [59]. Orphanage [71]. other [111]. Outline [97].

Paid [106]. Paper [71]. Papers [72, 63, 157]. Parentage [61]. Paris [111]. Parish [89, 155, 96, 137, 112]. Parishes [78, 149, 103]. Parliament [72, 63, 157]. Part [32]. Particulars [49]. Passengers [49]. Passing [106]. Pauper [32]. Pendleton [124]. Penzance [137]. Period [109]. Periods [57]. Physiology [121]. Platting [34]. Pointing [141]. Police [44, 136, 83]. Ponts [84]. Poor [42]. Poorest [59]. Population [89, 149, 87, 126, 143, 153, 57]. Port [110]. Porter [33]. Portion [95]. Potter [90]. Potteries [35]. Power [79]. Preceding [109, 126]. Present [106, 72, 63, 33, 131, 113, 157, 87, 57, 84]. Preston [125]. Previous [71]. Prices [106, 104]. Principal [61, 92, 57, 156]. Printed [72, 63, 157]. Prisoners [83]. Probable [61]. Proceedings [38, 51, 65, 105, 117, 123, 128, 134, 145]. Produce [57]. Progress [33, 131, 57, 84]. Provinces [127]. Provincial [37, 50]. Provisions [132]. Prussia [104, 131]. Prussian [131, 153]. Public [57]. Publication [36]. Publications [52, 66, 126]. Pulmonary [55].

Quantities [106].

R [33]. Railway [74, 49]. Railways [122]. Ramsbottom [112]. Rate [76]. Regents [90]. Registering [61]. Registrar [142]. Registrar-General [142]. Relations [33]. Relative [55]. Religious [89]. Remaining [106]. Remarks [126, 70]. Report [48, 78, 31, 90, 68, 74, 102, 130, 142, 152, 124, 150, 149, 136, 155, 96, 125, 111, 103, 154, 132, 34, 139]. Reports [84]. Result [61]. Retired [76]. Return [83]. Returns [43]. Rev. [125]. Revenue [33, 57]. Review [104, 109]. Roads [116]. Robberies [133]. Rutland [150]. Rutlandshire [149].

Saints [89]. Saving [111]. Saxony [48, 126]. Scholars [60]. School [78]. Schools [60, 32, 71]. Science [121, 82, 152, 147]. Scientific [156]. Seats [89]. Second [78, 36, 68, 74]. Section [114]. Sections [33]. Session [72, 63, 157]. Several [89]. Shewing [57]. Short [94]. Sickness [136, 139, 54, 69, 99]. Situation [95]. Six [126]. Social [33, 86]. Societies [37, 38, 50, 51, 65, 105, 117, 123, 128, 134, 145, 156]. Society [48, 78, 31, 36, 68, 102, 130, 141, 136, 126, 103, 124, 150]. Soho [103]. St. [78, 103]. State [42, 78, 90, 68, 102, 124, 150, 155, 34, 93]. Statement [89, 113, 57]. States [153, 90, 60]. Statistical [72, 63, 37, 38, 50, 51, 65, 105, 52, 66, 157, 117, 123, 128, 134, 145, 124, 150, 96, 137, 97, 92, 98, 139, 88, 84, 80, 75, 48, 78, 31, 68, 102, 130, 141, 136, 126, 103]. Statistics [152, 136, 86, 140, 126, 103, 143, 115, 114, 120, 40, 112, 148, 151, 85, 156]. Steam [79]. Steam-Power [79]. Stock [41]. Strikes [35]. Subjects [97]. Suffolk [62]. Suggestions [148]. Suicides [47]. Sunday [78]. Supplementary [103]. Systematic [148].

Table [98]. Tables [70]. Taxation [57]. Ten [83]. their [61]. Thereupon [70]. Third [102, 103]. those [126]. Three [149, 110]. Time [33]. Tin [140]. Tin-Mines [140]. Township [95, 124]. Trade [46, 131, 109, 110]. Training [32, 71]. Transmitting [90]. Trial [61]. Troops [139, 54, 69, 99]. Tulloch [139]. Turnpike [116]. Tyne [89, 152, 86, 88, 83].

Union [131, 32]. United [58, 90, 37, 50, 60, 113, 92, 139]. Universities [92]. University [90]. upon [89, 152, 136, 59, 86, 111, 88, 83].

Value [113, 121]. Various [33]. Vicinity [79]. View [84]. Vital [136]. Volume [29, 108, 119].

Wales [116, 151, 70]. Warehouse [106].

Wear [152]. Welsh [46]. West [155, 54, 69, 99]. Westminster [78, 68, 47, 102, 103]. Wheat [106]. Wheat-Flour [106]. which [89]. William [64]. within [89, 34]. Workhouses [32]. Working [155, 154].

Year [133, 150, 109, 111, 153]. Years [104, 35, 126, 110]. York [90, 120].

Zealand [87].

References

Anonymous:1834:N

[1] Anonymous. November 17, 1834. Proceedings of the Statistical Society of London, 1(1):1–13, November 1834. CO-DEN ???? ISSN 1466-6162. URL http:

                                             //www.jstor.org/stable/29765580.

Anonymous:1834:D

[2] Anonymous. December 15, 1834. Proceedings of the Statistical Society of London, 1(1):13–16, December 1834. CO-DEN ???? ISSN 1466-6162. URL http:

                                             //www.jstor.org/stable/29765581.

Anonymous:1835:Ja

[3] Anonymous. January 19, 1835. Proceedings of the Statistical Society of London, 1(2):9–11, January 1835. CODEN ???? ISSN 1466-6162. URL http://

                                             www.jstor.org/stable/29765584.

Anonymous:1835:F

[4] Anonymous. February 16, 1835. Proceedings of the Statistical Society of London, 1(2):12–17, February 1835. CO-DEN ???? ISSN 1466-6162. URL http:

//www.jstor.org/stable/29765585.

Anonymous:1835:Ma

[5] Anonymous. March 16, 1835. Proceedings of the Statistical Society of London, 1(2):17–25, March 1835. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/29765586.

Anonymous:1835:A

[6] Anonymous. April 20, 1835. Proceedings of the Statistical Society of London, 1(2):25–33, April 1835. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/29765587.

Anonymous:1835:Mb

[7] Anonymous. May 18, 1835. Proceedings of the Statistical Society of London, 1(2): 33–41, May 1835. CODEN ???? ISSN

1466-6162. URL http://www.jstor.
org/stable/29765588.

Anonymous:1835:Jb

[8] Anonymous. June 15, 1835. Proceedings of the Statistical Society of London, 1(2):41–52 54–59, June 1835. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/29765589.

Anonymous:1835:N

[9] Anonymous. November 16, 1835. Proceedings of the Statistical Society of London, 1(3):61–76, November 1835. CO-DEN ???? ISSN 1466-6162. URL http:

//www.jstor.org/stable/29765591.

Anonymous:1835:D

[10] Anonymous. December 21, 1835. Proceedings of the Statistical Society of London, 1(4):77–86, December 1835. CO-DEN ???? ISSN 1466-6162. URL http:

                                        //www.jstor.org/stable/29765593.

Hallam:1836:J

[11] Henry Hallam and W. H. Sykes. January 18, 1836. Proceedings of the Statistical Society of London, 1(4):87–98 100 102 104, January 1836. CODEN ???? ISSN

                                        1466-6162. URL http://www.jstor.
                                        org/stable/29765594.

Anonymous:1836:F

[12] Anonymous. February 15, 1836. Proceedings of the Statistical Society of London, 1(5):105–107, February 1836. CO-DEN ???? ISSN 1466-6162. URL http:

                                        //www.jstor.org/stable/29765596.

Anonymous:1836:Ma

[13] Anonymous. March 21, 1836. Proceedings of the Statistical Society of London, 1(5):108–118, March 1836. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/29765597.

Anonymous:1836:A

[14] Anonymous. April 18, 1836. Proceedings of the Statistical Society of London, 1(5):119–152, April 1836. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/29765598.

Anonymous:1836:Mb

[15] Anonymous. May 16, 1836. Proceedings of the Statistical Society of London, 1(6):153–168, May 1836. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/29765600.

Anonymous:1836:J

[16] Anonymous. June 20, 1836. Proceedings of the Statistical Society of London, 1(7):169–185, June 1836. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/29765602.

Anonymous:1836:N

[17] Anonymous. November 21, 1836. Proceedings of the Statistical Society of London, 1(8):187–191, November 1836. CO-DEN ???? ISSN 1466-6162. URL http:

//www.jstor.org/stable/29765604.

Anonymous:1836:D

[18] Anonymous. December 19, 1836. Proceedings of the Statistical Society of London, 1(8):192–208, December 1836. CO-DEN ???? ISSN 1466-6162. URL http:

//www.jstor.org/stable/29765605.

Anonymous:1837:J

[19] Anonymous. January 16, 1837. Proceedings of the Statistical Society of London, 1(8):209–221, January 1837. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/29765606.

Anonymous:1837:MF

[20] Anonymous. Monday, February 20, 1837. Proceedings of the Statistical Society of London, 1(9):223–235, February 1837. CODEN ???? ISSN 1466-6162.

URL http://www.jstor.org/stable/
29765608.

Anonymous:1837:MMa

[21] Anonymous. Monday, March 6, 1837. Proceedings of the Statistical Society of London, 1(9):236–260, March 1837. CO-DEN ???? ISSN 1466-6162. URL http:

//www.jstor.org/stable/29765609.

Anonymous:1837:M

[22] Anonymous. March 20, 1837. Proceedings of the Statistical Society of London, 1(10):261–267, March 1837. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/29765611.

Anonymous:1837:MA

[23] Anonymous. Monday, April 17, 1837. Proceedings of the Statistical Society of London, 1(10):268–285, April 1837. CO-DEN ???? ISSN 1466-6162. URL http:

//www.jstor.org/stable/29765612.

Anonymous:1837:MMb

[24] Anonymous. Monday, May 15, 1837. Proceedings of the Statistical Society of London, 1(10):286–294, May 1837. CO-DEN ???? ISSN 1466-6162. URL http:

//www.jstor.org/stable/29765613.

Anonymous:1837:MJ

[25] Anonymous. Monday, June 19, 1837. Proceedings of the Statistical Society of London, 1(11):295–308, June 1837. CO-DEN ???? ISSN 1466-6162. URL http:

//www.jstor.org/stable/29765615.

Anonymous:1837:MN

[26] Anonymous. Monday, November 20, 1837. Proceedings of the Statistical Society of London, 2(12):1–16, November 1837. CODEN ???? ISSN 1466-6162.

URL http://www.jstor.org/stable/
29765617.

Anonymous:1837:MD

[27] Anonymous. Monday, December 18, 1837. Proceedings of the Statistical Society of London, 2(13):17–24, December 1837. CODEN ???? ISSN 1466-6162.

URL http://www.jstor.org/stable/
29765619.

Anonymous:1838:MJ

[28] Anonymous. Monday, January 15, 1838. Proceedings of the Statistical Society of London, 2(14):25–31, January 1838. CODEN ???? ISSN 1466-6162.

URL http://www.jstor.org/stable/
29765621.

Anonymous:1838:VI

[29] Anonymous. Volume information. Journal of the Statistical Society of London, 1(1):i–iv, May 1838. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337848.

Anonymous:1838:I

[30] Anonymous. Introduction. Journal of the Statistical Society of London, 1(1): 1–5, May 1838. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337849.

Anonymous:1838:FAR

[31] Anonymous. Fourth annual report of the Council of the Statistical Society of London. Journal of the Statistical Society of London, 1(1):5–13, May 1838. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337850.

Kay:1838:ECD

[32] James Phillips Kay. On the establishment of county or district schools, for the training of the pauper children maintained in union workhouses. Part I. Journal of the Statistical Society of London, 1(1):14–27, May 1838. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337851.

Anonymous:1838:PNV

[33] Anonymous. The progress of the nation, in its various social and economical relations from the beginning of the Nineteenth Century to the present time. Sections III and IV: Interchange; and revenue and expenditure by G. R. Porter.

Journal of the Statistical Society of London, 1(1):27–33, May 1838. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337852.

Heywood:1838:REC

[34] James Heywood. Report of an enquiry, conducted from house to house, into the state of 176 families in miles platting, within the Borough of Manchester in 1837. Journal of the Statistical Society of London, 1(1):34–36, May 1838. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2337853.

Boyle:1838:ASP

[35] John Boyle. An account of strikes in the potteries, in the years 1834 and 1836. Journal of the Statistical Society of London, 1(1):37–45, May 1838. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337854.

Anonymous:1838:SPC

[36] Anonymous. Second publication by Central Society of Education. Journal of the Statistical Society of London, 1(1): 45–48, May 1838. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337855.

Anonymous:1838:PSSa

[37] Anonymous. Provincial statistical societies in the United Kingdom. Journal of the Statistical Society of London, 1(1): 48–50, May 1838. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337856.

Anonymous:1838:PSSb

[38] Anonymous. Proceedings of statistical societies. Journal of the Statistical Society of London, 1(1):50–52, May 1838. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337857.

Anonymous:1838:Ma

[39] Anonymous. Miscellaneaous. Journal of the Statistical Society of London, 1(1): 52–64, May 1838. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337858.

Lemon:1838:SCM

[40] Charles Lemon. The statistics of the copper mines of Cornwall. Journal of the Statistical Society of London, 1(2): 65–84, June 1838. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338061.

Anonymous:1838:JSC

[41] Anonymous. Joint-stock companies in France. Journal of the Statistical Society of London, 1(2):84–85, June 1838. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338062.

Anonymous:1838:AES

[42] Anonymous. An account of an enquiry into the state of 275 poor families in the city of Bristol. Journal of the Statistical Society of London, 1(2):86–88, June 1838. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338063.

Porter:1838:ARC

[43] G. R. Porter. Agricultural returns, County of Bedford. Journal of the Statistical Society of London, 1(2):89–96, June 1838. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338064.

Anonymous:1838:PM

[44] Anonymous. Police of the Metropolis. Journal of the Statistical Society of London, 1(2):96–103, June 1838. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338065.

Phillips:1838:MA

[45] Benjamin Phillips. Mortality of amputation. Journal of the Statistical Society of London, 1(2):103–105, June 1838. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338066.

Anonymous:1838:WFT

[46] Anonymous. Welsh flannel trade. Journal of the Statistical Society of London, 1(2):105–107, June 1838. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338067.

Anonymous:1838:SW

[47] Anonymous. Suicides in Westminster from 1812–1836. Journal of the Statistical Society of London, 1(2):107–110, June 1838. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338068.

Anonymous:1838:ARS

[48] Anonymous. Annual report of the Statistical Society of Saxony. Journal of the Statistical Society of London, 1(2):110– 114, June 1838. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338069.

Loch:1838:PCC

[49] George Loch. Particulars connected with the conveyance of passengers on the Brussels and Antwerp railway. Journal of the Statistical Society of London, 1(2):

114–115, June 1838. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338070.

Anonymous:1838:PSSc

[50] Anonymous. Provincial statistical societies in the United Kingdom. Journal of the Statistical Society of London, 1(2): 115–117, June 1838. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338071.

Anonymous:1838:PSSd

[51] Anonymous. Proceedings of statistical societies. Journal of the Statistical Society of London, 1(2):117–118, June 1838. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338072.

Anonymous:1838:RSPa

[52] Anonymous. Recent statistical publications. Journal of the Statistical Society of London, 1(2):118–120, June 1838. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338073.

Anonymous:1838:Mb

[53] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 1(2): 121–128, June 1838. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338074.

Tulloch:1838:SMAa

[54] A. M. Tulloch. On the sickness and mortality among the troops in the West Indies. Journal of the Statistical Society of London, 1(3):129–142, July 1838. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2337906.

Clendinning:1838:RFP

[55] John Clendinning. On the relative frequency of pulmonary consumption and diseases of the heart. Journal of the Statistical Society of London, 1(3):142– 149, July 1838. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337907.

Anonymous:1838:ECC

[56] Anonymous. Endowed charities in Cornwall. Journal of the Statistical Society of London, 1(3):149–153, July 1838. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2337908.

Porter:1838:SPP

[57] G. R. Porter. Statement of the progress of the principal branches of the public revenue, at different periods during the present century, shewing the actual produce of taxation, and its produce in comparison with the increased population. Journal of the Statistical Society of London, 1(3):154–155, July 1838. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337909.

Anonymous:1838:EUK

[58] Anonymous. Emigration from the United Kingdom. Journal of the Statistical Society of London, 1(3):155–167, July 1838. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2337910.

Baird:1838:OUP

[59] C. R. Baird. Observations upon the poorest class of operatives in Glasgow in 1837. Journal of the Statistical Society of London, 1(3):167–172, July 1838. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2337911.

Anonymous:1838:SSM

[60] Anonymous. Schools and scholars in Massachusetts, United States, 1837. Journal of the Statistical Society of London, 1(3):173–174, July 1838. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337912.

Anonymous:1838:FRP

[61] Anonymous. Forms for registering the principal circumstances connected with the birth, parentage, education, and condition of criminal offenders; the nature and probable causes of their offences, with the result of their trial. Journal of the Statistical Society of London, 1(3):174–178, July 1838. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337913.

Phillips:1838:EAL

[62] James Phillips. Earnings of agricultural labourers in Norfolk and Suffolk. Journal of the Statistical Society of London, 1(3):179–183, July 1838. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337914.

Anonymous:1838:LSP

[63] Anonymous. A list of statistical papers printed by the Houses of Parliament during the present session of 1837– 8. Journal of the Statistical Society of London, 1(3):183–186, July 1838. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2337915.

Anonymous:1838:AAE

[64] Anonymous. An abstract of the account of the expenses of the coronations of George the fourth and of William the fourth. Journal of the Statistical Society of London, 1(3):187–189, July 1838. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2337916.

Anonymous:1838:PSSe

[65] Anonymous. Proceedings of statistical societies. Journal of the Statistical Society of London, 1(3):189–190, July 1838. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337917.

Anonymous:1838:RSPb

[66] Anonymous. Recent statistical publications. Journal of the Statistical Society of London, 1(3):190–191, July 1838. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2337918.

Anonymous:1838:Mc

[67] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 1(3): 191–192, July 1838. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337919.

Anonymous:1838:SRC

[68] Anonymous. Second report of a committee of the Statistical Society of London, appointed to enquire into the state of education in Westminster. Journal of the Statistical Society of London, 1(4):193– 215, August 1838. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337801.

Tulloch:1838:SMAb

[69] A. M. Tulloch. On the sickness and mortality among the troops in the West Indies. Journal of the Statistical Society of London, 1(4):216–230, August 1838. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337802.

Redgrave:1838:ACT

[70] S. Redgrave. Abstract of criminal tables for England and Wales, 1837; with remarks thereupon, chiefly by the compiler. Journal of the Statistical Society of London, 1(4):231–245, August 1838. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337803.

Kay:1838:NIP

[71] James Phillips Kay. Notes illustrative of a previous paper on the training in schools of industry of children dependent from crime, orphanage. Journal of the Statistical Society of London, 1(4):245– 251, August 1838. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337804.

Anonymous:1838:FLS

[72] Anonymous. A further list of statistical papers printed by the Houses of Parliament during the present session of 1837–8. Journal of the Statistical Society of London, 1(4):251–252, August 1838. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337805.

Anonymous:1838:Md

[73] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 1(4): 253–256, August 1838. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2337806.

Anonymous:1838:SRI

[74] Anonymous. Second report of the Irish Railway Commissioners. Journal of the Statistical Society of London, 1(5):257– 277, September 1838. CODEN ????

ISSN 0959-5341. URL http://www.

jstor.org/stable/2337983.

Thomson:1838:SEF

[75] Arthur Saunders Thomson. A statistical enquiry on fever. Journal of the Statistical Society of London, 1(5):278–279, September 1838. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337984.

Christie:1838:RMA

[76] Robert Christie. On the rate of mortality amongst officers retired from the Indian Army. Journal of the Statistical Society of London, 1(5):279–283, September 1838. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337985.

Anonymous:1838:FLA

[77] Anonymous. Fires in London, with an account of the London fire-engine establishment. Journal of the Statistical Society of London, 1(5):283–297, September 1838. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337986.

Anonymous:1838:ASR

[78] Anonymous. Appendix to the second report of the Statistical Society of London on the state of education in Westminster, being a detailed account of each endowed or charity school, and of each infant and Sunday school, in the parishes of St. Margaret and St. John, Westminster. Journal of the Statistical Society of London, 1(5):298–315, September 1838. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337987.

Anonymous:1838:ISP

[79] Anonymous. Increase of steam-power in Lancashire and its immediate vicinity. Journal of the Statistical Society of London, 1(5):315–316, September 1838. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2337988.

Portlock:1838:AEO

[80] J. E. Portlock. An address explanatory of the objects and advantages of statistical enquiries. Journal of the Statistical Society of London, 1(5):316–317, September 1838. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337989.

Anonymous:1838:Me

[81] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 1 (5):317–320, September 1838. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337990.

Anonymous:1838:EMB

[82] Anonymous. Eighth meeting of the British Association for the Advancement of Science. Journal of the Statistical Society of London, 1(6):321–324, October 1838. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337749.

Stephens:1838:ARP

[83] John Stephens. Abstract of a return of prisoners coming under the cognizance of the police of Newcastle-upon-Tyne, during the ten months from the 2nd of October, 1837 to the 2nd of August, 1838. Journal of the Statistical Society of London, 1(6):324–326, October 1838. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337750.

Porter:1838:SVR

[84] G. R. Porter. A statistical view of the recent progress and present amount of mining industry in France; drawn from the official reports of the “Direction Générale des Ponts et Chaussées et des Mines”. Journal of the Statistical Society of London, 1(6):326–336, October 1838. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337751.

Sykes:1838:SC

[85] W. H. Sykes. Statistics of Cadiz. Journal of the Statistical Society of London, 1(6):337–354, October 1838. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337752.

Cargill:1838:ECS

[86] William Cargill. Educational, criminal, and social statistics of Newcastle-upon-Tyne. Journal of the Statistical Society of London, 1(6):355–361, October 1838. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337753.

Bannister:1838:ACP

[87] Saxe Bannister. An account of the changes and present condition of the population of New Zealand. Journal of the Statistical Society of London, 1 (6):362–376, October 1838. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337754.

MAlister:1838:SNA

[88] Joseph M’Alister. Statistical notice of the asylum for the blind in Newcastleupon-Tyne. Journal of the Statistical Society of London, 1(6):377–379, October 1838. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2337755.

Anonymous:1838:SNS

[89] Anonymous. A statement of the number of seats in the several churches and chapels, distinguishing the several religious denominations to which they belong, within the Parish of All Saints, Newcastle-upon-Tyne, containing a population of about 20,000 inhabitants, in the month of August, 1838. Journal of the Statistical Society of London, 1 (6):379–382, October 1838. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337756.

Anonymous:1838:LDP

[90] Anonymous. A letter from Dr. Potter, of the United States, transmitting the last annual report of the Regents of the University of the State of New York. Journal of the Statistical Society of London, 1(6):383–384, October 1838. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337757.

Anonymous:1838:Mf

[91] Anonymous. Miscellaneaous. Journal of the Statistical Society of London, 1(6): 384, October 1838. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337758.

Jones:1838:SIP

[92] H. Longueville Jones. Statistical illustrations of the principal universities of the United Kingdom of Great Britain and Ireland. Journal of the Statistical Society of London, 1(7):385–397, November 1838. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2337769.

Hindmarsh:1838:SAC

[93] L. Hindmarsh. On the state of agriculture and condition of the agricultural labourers of the Northern Division of Northumberland. Journal of the Statistical Society of London, 1(7):397–414, November 1838. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337770.

Wilson:1838:SAD

[94] Thos. Wilson. A short account of the Darton Collieries Club. Journal of the Statistical Society of London, 1(7):414– 416, November 1838. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2337771.

Felkin:1838:ASP

[95] W. Felkin. An account of the situation of a portion of the labouring classes in the township of Hyde, Cheshire. Journal of the Statistical Society of London, 1(7):416–420, November 1838. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337772.

Charlton:1838:SRP

[96] W. H. Charlton. A statistical report from the Parish of Bellingham in the County of Northumberland. Journal of the Statistical Society of London, 1 (7):420–426, November 1838. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2337773.

Hare:1838:AOS

[97] Samuel Hare. Abstract of outline of subjects for statistical enquiries. Journal of the Statistical Society of London, 1 (7):426–427, November 1838. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337774.

Kingsley:1838:STC

[98] J. Kingsley. A statistical table of crime in Ireland. Journal of the Statistical Society of London, 1(7):427, November 1838. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337775.

Tulloch:1838:SMAc

[99] A. M. Tulloch. On the sickness and mortality among the troops in the West Indies. Journal of the Statistical Society of London, 1(7):428–444, November 1838. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337776.

Anonymous:1838:NE

[100] Anonymous. Note by the Editor. Journal of the Statistical Society of London, 1(7):444–445, November 1838. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337777.

Anonymous:1838:Mg

[101] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 1 (7):445–448, November 1838. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337778.

Anonymous:1838:TRC

[102] Anonymous. Third report of a committee of the Statistical Society of London appointed to enquire into the state of education in Westminster. Journal of the Statistical Society of London, 1 (8):449–477, December 1838. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338093.

Edgell:1838:MSP

[103] Edgell Wyatt Edgell. Moral statistics of the parishes of St. James, St. George, and St. Anne Soho, in the City of Westminster. Supplementary to the third report of the Education Committee of the Statistical Society of London. Journal of the Statistical Society of London, 1 (8):478–492, December 1838. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338094.

Anonymous:1838:RPC

[104] Anonymous. A review of the prices of corn in Prussia during the 22 years from 1816 to 1837. Journal of the Statistical Society of London, 1(8):493–505, December 1838. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338095.

Anonymous:1838:PSSf

[105] Anonymous. Proceedings of statistical societies. Journal of the Statistical Society of London, 1(8):506–507, December 1838. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338096.

Anonymous:1838:AQF

[106] Anonymous. An account of the quantities of foreign and colonial wheat, and wheat-flour, imported, paid duty, and remaining in warehouse; and the monthly average prices of British wheat, in each month from the passing of the Act 6 Geo. IV. c. 60, on the 15th July, 1828, to the present date. Journal of the Statistical Society of London, 1(8): 507–511, December 1838. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338097.

Anonymous:1838:Mh

[107] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 1 (8):512, December 1838. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338098.

Anonymous:1839:VIa

[108] Anonymous. Volume information. Journal of the Statistical Society of London, 1(9):559–572, January 1839. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338075.

Anonymous:1839:RTF

[109] Anonymous. A review of the trade of France in the year 1837, compared with the preceding year and with the decennial period from 1827 to 1836. Journal of the Statistical Society of London, 1(9):513–522, January 1839. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338076.

Sturrock:1839:ATP

[110] John Sturrock and Joseph Hume. An account of the trade of the Port of Dundee, during the three years ended 31st May, 1838. Journal of the Statistical Society of London, 1(9):522–529, January 1839. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338077.

Delessert:1839:RUO

[111] M. Benjamin Delessert and H. Longueville Jones. Report upon the operations of the Paris and other savings’ banks in France, during the year 1837. Journal of the Statistical Society of London, 1(9):529–537, January 1839. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338078.

MDouall:1839:SPR

[112] P. M. M’Douall. Statistics of the Parish of Ramsbottom, near Bury, in Lancashire. Journal of the Statistical Society of London, 1(9):537–539, January 1839. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338079.

Anonymous:1839:ASP

[113] Anonymous. An approximate statement of the present extent and value of the manufacture of glass in the United Kingdom. Journal of the Statistical Society of London, 1(9):539–540, January 1839. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338080.

Felkin:1839:MSS

[114] W. Felkin. Moral statistics of a section of the City of Norwich. Journal of the Statistical Society of London, 1(9):540– 541, January 1839. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338081.

Felkin:1839:MSD

[115] W. Felkin. Moral statistics of a district near Gray’s inn, London, in 1836. Journal of the Statistical Society of London, 1(9):541–542, January 1839. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338082.

Anonymous:1839:TRE

[116] Anonymous. Turnpike roads in England and Wales. Journal of the Statistical Society of London, 1(9):542–545, January 1839. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338083.

Anonymous:1839:PSSa

[117] Anonymous. Proceedings of statistical societies. Journal of the Statistical Society of London, 1(9):546–554, January 1839. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338084.

Anonymous:1839:Ma

[118] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 1 (9):555–558, January 1839. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338085.

Anonymous:1839:VIb

[119] Anonymous. Volume information. Journal of the Statistical Society of London, 2(1):i–iv, February 1839. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337807.

Fripp:1839:SCN

[120] Charles Bowles Fripp. Statistics of the city of new York. Journal of the Statistical Society of London, 2(1):1–25, February 1839. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2337808.

Guy:1839:VNM

[121] William Augustus Guy. On the value of the numerical method as applied to science, but especially to physiology and medicine. Journal of the Statistical Society of London, 2(1):25–47, February 1839. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337809.

Anonymous:1839:RB

[122] Anonymous. On railways in Belgium.

Journal of the Statistical Society of London, 2(1):47–62, February 1839. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2337810.

Anonymous:1839:PSSb

[123] Anonymous. Proceedings of statistical societies. Journal of the Statistical Society of London, 2(1):63–64, February 1839. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337811.

Anonymous:1839:RCMa

[124] Anonymous. Report of a Committee of the Manchester Statistical Society, on the state of education in the Township of Pendleton, 1838. Journal of the Statistical Society of London, 2(2):65–83, March 1839. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337977.

Clay:1839:ARR

[125] John Clay. Annual report of the Rev. John Clay, Chaplain to the Preston House of Correction. Journal of the Statistical Society of London, 2(2):84– 103, March 1839. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337978.

Deverell:1839:SPK

[126] W. R. Deverell. Statistics of the population of the Kingdom of Saxony, to December 1837, including an account of the births, deaths, and marriages, compared with those of the six preceding years. Abstracted from the publications of the Statistical Society of Saxony. With incidental and general remarks. Journal of the Statistical Society of London, 2 (2):103–114, March 1839. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2337979.

Anonymous:1839:AAF

[127] Anonymous. An account of Algeria, or the French provinces in Africa. Journal of the Statistical Society of London, 2(2):115–126, March 1839. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337980.

Anonymous:1839:PSSc

[128] Anonymous. Proceedings of statistical societies. Journal of the Statistical Society of London, 2(2):126, March 1839. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337981.

Anonymous:1839:Mb

[129] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 2(2): 127–128, March 1839. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2337982.

Anonymous:1839:FARa

[130] Anonymous. Fifth annual report of the Council of the Statistical Society of London. Journal of the Statistical Society of London, 2(3):129–134, April 1839. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338019.

Anonymous:1839:ARP

[131] Anonymous. An account of the recent progress and present extent of manufactures in Prussia, and of the trade of the Prussian commercial union in manufactured goods. Journal of the Statistical Society of London, 2(3):135–172, April 1839. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338020.

Fry:1839:RIF

[132] Alfred A. Fry. Report of the inspectors of factories on the effects of the educational provisions of the Factories’ Act. Journal of the Statistical Society of London, 2(3):173–181, April 1839. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338021.

Anonymous:1839:NNR

[133] Anonymous. Number and nature of robberies in London and Liverpool during the year 1838. Journal of the Statistical Society of London, 2(3):182–184, April 1839. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338022.

Anonymous:1839:PSSd

[134] Anonymous. Proceedings of statistical societies. Journal of the Statistical Society of London, 2(3):185–188, April 1839. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338023.

Anonymous:1839:Mc

[135] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 2 (3):188–192, April 1839. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338024.

Anonymous:1839:RCS

[136] Anonymous. Report of a committee of the Statistical Society of London, appointed to collect and enquire into vital statistics, upon the sickness and mortality among the Metropolitan Police Force. May, 1839. Journal of the Statistical Society of London, 2(4):193–197, July 1839. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2337706.

Edmonds:1839:SAP

[137] Richard Edmonds. A statistical account of the Parish of Madron, containing the Borough of Penzance, in Cornwall. Journal of the Statistical Society of London, 2(4):198–233, July 1839. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337707.

Whishaw:1839:AEC

[138] James Whishaw. An account of the endowed charities in Herefordshire. Journal of the Statistical Society of London, 2(4):234–250, July 1839. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337708.

Lever:1839:SMA

[139] J. W. C. Lever. On the sickness and mortality among the troops in the United Kingdom. Abstract of the statistical report of Major Tulloch. Journal of the Statistical Society of London, 2(4):250– 260, July 1839. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337709.

Carne:1839:STM

[140] Joseph Carne. Statistics of the tin-mines in Cornwall, and of the consumption of tin in Great Britain. Journal of the Statistical Society of London, 2(4):260– 268, July 1839. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337710.

Anonymous:1839:LDJ

[141] Anonymous. Letter from Dr. Julius, foreign member of the Statistical Society of London, &c., pointing out the excess of male births among Jews in the city of Hamburg. Journal of the Statistical Society of London, 2(4):268, July 1839. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337711.

Anonymous:1839:FARb

[142] Anonymous. First annual report of the Registrar-General on births, deaths, and marriages in England, in 1837–8. Journal of the Statistical Society of London, 2(4):269–274, July 1839. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337712.

Fallati:1839:SPN

[143] Dr. Fallati. Statistics of the population of Norway. Journal of the Statistical Society of London, 2(4):275–278, July 1839. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337713.

David:1839:NLC

[144] C. N. David. A notice on the last census in Denmark. Journal of the Statistical Society of London, 2(4):279–280, July 1839. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337714.

Anonymous:1839:PSSe

[145] Anonymous. Proceedings of statistical societies. Journal of the Statistical Society of London, 2(4):281–285, July 1839. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337715.

Anonymous:1839:Md

[146] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 2(4):

285–288, July 1839. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337716.

Anonymous:1839:NMB

[147] Anonymous. Ninth meeting of the British Association for the Advancement of Science — August, 1839. Journal of the Statistical Society of London, 2 (5):289–290, October 1839. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337817.

Porter:1839:SFS

[148] G. R. Porter. Suggestions in favour of the systematic collection of the statistics of agriculture. Journal of the Statistical Society of London, 2(5):291–296, October 1839. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337818.

Anonymous:1839:RCP

[149] Anonymous. Report on the condition of the population in three parishes in Rutlandshire, in March, 1839. Journal of the Statistical Society of London, 2(5):297– 302, October 1839. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337819.

Anonymous:1839:RCMb

[150] Anonymous. Report of a Committee of the Manchester Statistical Society on the state of education in the County of Rutland in the year 1838. Journal of the Statistical Society of London, 2(5):303– 315, October 1839. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337820.

Rawson:1839:ISC

[151] Rawson W. Rawson. An inquiry into the statistics of crime in England and Wales. Journal of the Statistical Society of London, 2(5):316–344, October 1839. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2337821.

Anonymous:1839:FRC

[152] Anonymous. First report of a committee of the British Association for the Advancement of Science, appointed at Newcastle, to inquire into the statistics of the collieries upon the Tyne and Wear. Journal of the Statistical Society of London, 2(5):345–356, October 1839. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2337822.

Hebeler:1839:CPP

[153] Bernard Hebeler. Changes in the population of the Prussian states, during the year 1838. Journal of the Statistical Society of London, 2(5):356–368, October 1839. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2337823.

Fripp:1839:RIC

[154] C. Bowles Fripp. Report of an inquiry into the condition of the working classes of the city of Bristol. Journal of the Statistical Society of London, 2(5):368– 375, October 1839. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2337824.

Anonymous:1839:RSE

[155] Anonymous. Report on the state of education among the working classes in the Parish of West Bromwich. Journal of the Statistical Society of London, 2(5):375– 377, October 1839. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2337825.

Whishaw:1839:SPS

[156] James Whishaw. Statistics of the principal scientific and literary societies in London. Journal of the Statistical Society of London, 2(5):378, October 1839. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2337826.

Anonymous:1839:FLS

[157] Anonymous. A further list of statistical papers printed by the Houses of Parliament during the present session of 1839. Journal of the Statistical Society of London, 2(5):378–380, October 1839. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2337827.

Anonymous:1839:Me

[158] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 2 (5):381–384, October 1839. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2337828.